OpenStudy (ineedhelpfast12):

Solve the linear system using elimination. x+y=2 2x+7y=9

OpenStudy (solomonzelman):

you can multiply the first equation times -2, and add the equations.

OpenStudy (ineedhelpfast12):

so -2x-2y=-4 2x=7y=9

OpenStudy (solomonzelman):

yes, correct -2x -2y = -4 2x + 7y = 9

OpenStudy (solomonzelman):

Now, add them.

OpenStudy (ineedhelpfast12):

y=1

OpenStudy (solomonzelman):

Yes, that was quick!

OpenStudy (solomonzelman):

And then plug y=1 into one of the original equations

OpenStudy (solomonzelman):

and solve for x.

OpenStudy (ineedhelpfast12):

x=1

OpenStudy (solomonzelman):

Yes, fabulous!

OpenStudy (solomonzelman):

So your solution is: x=1 and y=1 or, that could be recorded as a point, (1,1)

OpenStudy (ineedhelpfast12):

Thank you! Could you help me with a couple or so more?

OpenStudy (solomonzelman):

Hopefully I will have time. Will see how it goes:)

OpenStudy (ineedhelpfast12):

3x-2y=3 -x+y=1